# Consumption Partial Insurance

### v4

@@ Abstract
We measure consumption insurance against income shocks accounting for high-order moments of the income distribution. We derive a nonlinear consumption function, where the extent of insurance varies with the shocks’ sign and magnitude. Using PSID data, we estimate an asymmetric pass-through of bad versus good permanent shocks – 17% of a $3\sigma$ negative shock transmits to consumption versus 9% of an equal-sized positive shock – with greater pass-through as shocks worsen. Households would sacrifice over 1/8 of lifetime consumption to eliminate tail income risk. Our results align with survey responses to hypothetical events and suggest that tail risk matters substantially for consumption.
@@ 1 Introduction
How does consumption respond to income shocks? The answer to this question is important for determining the welfare effects of shifts in the income distribution, i.e., the link between income and consumption inequality, for understanding how families cope after adverse events, for designing government insurance, and for the dynamics of business cycles. A consistent empirical finding is that consumption is partially insured against income shocks, so income fluctuations translate into consumption fluctuations less than one-to-one (e.g., Blundell et al., 2008; Heathcote et al., 2014). [^1] The study of consumption insurance dates back to tests of the permanent income (Deaton and Paxson, 1994) and complete markets hypotheses (Cochrane, 1991; Attanasio and Davis, 1996; Hayashi et al., 1996; Brav et al., 2002). Jappelli and Pistaferri (2010) and Crawley and Theloudis (2024) review this literature. With a few exceptions discussed subsequently, most measures of consumption insurance do not distinguish between good and bad shocks, or between small and large ones. However, recent evidence shows that the distribution of income risk has a long and fat left tail; that is, there is an asymmetrically large probability that households face some really bad income shocks. This paper measures the consumption response to income shocks precisely when the distribution of said shocks may exhibit long and fat tails.
@@
If most income shocks away from the mean are large and bad shocks, their effect on consumption might, intuitively, differ from that of small/moderate shocks. In other words, left skewness and excess kurtosis, which are statements about the third and fourth moments of the income distribution respectively, may matter for the pass-through of shocks into consumption. This is precisely what we investigate in this paper. Specifically, we measure the degree of consumption insurance to income shocks of varying persistence, accounting for higher-order moments of their distribution (i.e., tail shocks). This is in contrast to the methodology in BPP, which revolves around the second moments of the joint income-consumption distribution and, thus, abstracts from the modern higher-order facts that characterize income dynamics (the second moment is a poor summary of the tails).
We offer four main contributions:
@@
First, we re-estimate BPP using recent PSID data over 1999–2019, targeting the original income and consumption second moments. However, while BPP imputed consumption from the CEX (the PSID previously lacked comprehensive consumption data), such data are now internally available in the PSID. We subsequently use this data to study tail income risk, so it is natural to start with a ‘replication’ of BPP using the newer data. This ‘replication’ is insightful in its own right not only because of the different time period, but also because we no longer need to rely on imputed consumption from the CEX.
@@
Second, we introduce into BPP the third and fourth moments of the joint distribution of income and consumption, in addition to the second moments that BPP originally use. This is our first attempt to go beyond the workhorse income-consumption covariance and measure partial insurance targeting additional information in the tails of the distribution.
@@
Third, we generalize the method of measuring consumption insurance to explicitly permit a role for tail income risk. BPP log-linearize the policy rule of a lifecycle consumption/savings problem to obtain a consumption function that depends linearly on income shocks. Our first and second contributions above revolve firmly around such a linear framework. We argue (as also does Carroll, 2001) that linearity is a poor approximation to behavior when income is subject to tail shocks. We then derive a higher-order (quadratic) consumption function that allows for nonlinear transmission of shocks. We characterize and empirically measure the consumption pass-through of income shocks, which, in this environment, depends explicitly on the magnitude and sign of shocks. While a nonlinear consumption function has been estimated before, most notably in Arellano et al. (2017) – henceforth ABB, our framework is far less data demanding (we need second moments of consumption and up to fourth moments of income, rather than observing the entire consumption-income joint distribution) and offers a more transparent link with the higher-order features of the distribution of income shocks. Both points have certain important advantages, which we discuss below.
Fourth, we estimate the welfare costs of tail income risk by calculating what fraction of lifetime consumption a household is willing to give up to avoid tail risk. We use parametric utility to calculate these costs, in line with De Nardi et al. (2019), Busch and Ludwig (2024), and Guvenen et al. (2024), but we base these calculations on the pass-through measure from the quadratic consumption function of the previous step. This measures an ‘aggregate’ of all partial insurance available to households, so it is not specific to a particular structural mechanism or a particular parameterization of preferences and expectations.
@@
Our baseline empirical implementation uses income and consumption panel data from the PSID over the years 1999–2019. We focus on a representative sample of stable married couples that is similar to the sample in BPP. We show that unexplained income growth in this sample features large negative skewness and excess kurtosis, qualitatively and numerically close to the skewness and kurtosis in the administrative data of Guvenen et al. (2021).
We then obtain five main findings:
@@
First, permanent and transitory shocks to disposable household income are negatively skewed, with skewness coefficients at $-0.95$ and $-1.31$, respectively, and highly leptokurtic, with kurtosis coefficients at $44.39$ and $49.23$, respectively. Left skewness implies that a negative shock is typically more unsettling, more hurtful than a positive shock, because it is further away from the zero mean. Excess kurtosis implies that more shocks are concentrated either in the middle of the distribution or far out in the tails, relative to the shoulders. Taken together, large left skewness and excess kurtosis suggest that the realizations of income shocks tend to be either close to their zero mean or far out in the left tail; that is, income shocks for most households are either small/average or quite bad. However, these features vary by age and by the household’s position in the income distribution.
@@
Second, in estimating BPP in the recent years, we find full insurance to transitory shocks and a small pass-through of permanent shocks at 0.15 versus 0.64 in BPP. A 10% permanent income cut thus reduces consumption by only 1.5%, vis-à-vis 6.4% in BPP. In other words, we find much larger partial insurance to permanent shocks in the recent years than originally found in 1980–1992.222Using recent PSID data but a different method, ABB and Arellano et al. (2023) find a similarly low pass-through of permanent shocks. We discuss this further when we present our empirical results. We attribute about a third of the difference from BPP to the consumption imputation from the CEX that BPP use. The imputation brings in large amounts of measurement error, inflating the covariance between income and consumption growth. Another third of the difference arises because the modern PSID provides data biennially (versus annually previously), which tends to mute the transmission of higher frequency shocks. The last third reflects the different calendar times.
@@
Third, introducing the third and fourth moments of income and consumption growth into BPP’s linear consumption setting does not change our estimates of partial insurance. This is counterintuitive in light of the literature that highlights, in multiple settings, the large welfare costs of tail risk (e.g., Barro, 2009; De Nardi et al., 2019). One would thus expect that targeting higher-order moments would reduce the degree of partial insurance, which is not what we find. This result is suggestive of a potential misspecification in the linear consumption function, which motivates our generalization of the methodology through the introduction of a higher-order consumption specification subsequently.
@@
Fourth, in estimating our quadratic consumption function, we find an asymmetry in the pass-through of bad versus good shocks, and of small versus large ones. In this environment, the pass-through depends on the magnitude and sign of shocks; and targeting higher-order income moments enables the identification of the transmission parameters. We draw three main conclusions: 1.) Bad permanent shocks are more unsettling than equal-sized good ones because they have larger transmission rates into consumption – a three standard deviations negative shock (a 50% permanent income cut) has a pass-through of $0.17$ (17% of the shock transmits into consumption), while a positive shock of similar magnitude has a pass-through of $0.09$. 2.) The pass-through of bad shocks increases with the severity of the shock. Bad permanent shocks are thus double hurtful: not only are they more extreme than good ones (they lie further out in the left tail), but they also transmit increasingly more as their magnitude worsens. 3.) Transitory shocks are fully insured in the full sample, at least from a statistical point of view, although the point estimates of their transmission parameters suggest that large transitory windfalls shift consumption considerably, consistent with quasi-experimental evidence of large responses to transitory gains (e.g., Fagereng et al., 2021). Among low wealth low educated households, however, transitory shocks, and in particular negative ones, shift consumption significantly. Overall, the results are consistent with an environment in which households save a good income shock but dissave in the presence of a bad one. There is clearly a limit to how much one can dissave or borrow, so a large negative shock ends up affecting consumption by relatively more.
@@
The asymmetric transmission of bad versus good permanent shocks varies with age and the household’s position in the income distribution. While older households insure average permanent shocks better than younger ones, older households have larger transmission rates of extreme shocks. Severely bad permanent shocks thus cause more harm among the old. Moreover, bad permanent shocks from a given position in the bottom tail of the distribution of income risk cause more harm among the rich (households in the top of the income distribution) than among the poor (households in the bottom); the converse is true for good shocks. Guvenen et al. (2021) show that income skewness becomes more negative with age and income level – there is more room for income to fall among older or higher income workers. In turn, we document that there is also more room for consumption to fall. Among the old, this is because both the rate of transmission and the severity of downwards income risk increase with age; among the rich, however, the asymmetric transmission is driven mostly by the increasing severity of downwards risk along the income distribution.
[…]
@@
There is a large literature in macro and labor economics that studies the transmission of income shocks into consumption, i.e., the link between income and consumption inequality. There are two distinct strands in this literature: a quasi-reduced form branch that measures the degree of partial insurance in the data without fully specifying the underlying insurance mechanisms, and a structural branch that fully models specific insurance channels.
@@
The seminal paper in the first branch, namely BPP, sparked a large volume of research. [^4] Hall and Mishkin (1982), Deaton and Paxson (1994) and Blundell and Preston (1998) characterize the empirical joint distribution of consumption and income, so they are early precursors in this literature. Blundell et al. (2016) measure partial insurance against wage risk, allowing for family labor supply. Theloudis (2021) allows for preference heterogeneity in partial insurance, which he estimates using information in the third moments. Commault (2022) allows consumption to respond to past transitory shocks and finds a strong response, similar to what experimental studies find. Hryshko and Manovskii (2022) identify sets of households in the PSID with different degrees of insurance based on certain features of their incomes. Crawley and Kuchler (2023) address neglected time aggregation in income and consumption in the PSID.
The previous papers restrict income shocks to transmit linearly into consumption. ABB and Arellano et al. (2023) generalize the income and consumption processes by letting income feature nonlinear persistence, which varies with the level of income and its shock, and nonlinear transmission. Using the recent PSID and a quantile regression approach, they measure the marginal propensity to consume from persistent income, which varies along the income distribution. They simulate the transmission of shocks and document asymmetries between bad/good shocks or between rich/poor, similar to ours. Our papers, however, differ in important dimensions. First, identification in ABB is quite demanding as they require observing the entire consumption-income joint distribution, while we need only second moments of consumption and up to fourth moments of income. This is appealing because we use more accessible information and because, unlike income, we cannot assess the higher-order features of consumption against an established population benchmark. The price of our econometrically more accessible approach is that, unlike ABB, we have to maintain a linear income process (but subject to tail shocks). Yet, we find that nonlinear persistence is not crucial for the main consumption results, i.e., the asymmetric impact of different shocks or at different income levels. Second, ABB measure the pass-through over quantiles of the distribution of income shocks, without explicitly studying the higher-order features of their distribution. We show subsequently that seemingly different quantiles, e.g., shocks in the 10th or 90th percentile as in ABB, may in fact be very similar to one another when the underlying distribution is skewed and leptokurtic. Our model allows us to estimate pass-through rates that explicitly shift with the sign and size of income shocks and can thus describe more transparently how shocks of different signs/sizes shift consumption.
@@
The asymmetric response to bad and good shocks is recently also found in surveys that elicit responses to hypothetical events. Bunn et al. (2018) find that the marginal propensity to consume from bad shocks is larger than from good shocks; Christelis et al. (2019) and Fuster et al. (2020) find that responses to losses are larger than to gains, and the response increases with the severity of the loss. We document that these patterns exist also in survey data, where, unlike elicitation surveys, income shocks are realized but unobserved. [^6] Given our finding of heterogeneous responses to shocks, our paper is also broadly related to the literature that studies MPC heterogeneity (e.g., Lewis et al., 2022; Crawley and Kuchler, 2023).
The paper is finally related to the newer income dynamics literature that focuses on non-Gaussian aspects. Geweke and Keane (2000) and Bonhomme and Robin (2010) use the PSID to document, respectively, left skewness and excess kurtosis in earnings growth. Guvenen et al. (2021) use administrative data in the U.S. to provide further evidence of left skewness and excess kurtosis in earnings growth, accentuated with age and the level of earnings. Guvenen et al. (2014) and Busch et al. (2018) study the cyclical nature of skewness in several countries. While these papers are mostly descriptive of the higher-order features of income, we further document the implications of these features for consumption. We thus contribute to a growing literature on the implications of tail risk, e.g., for growth (Barro, 2009), asset prices (Constantinides and Ghosh, 2017), business cycles (McKay, 2017), consumer welfare (De Nardi et al., 2019), and labor market dynamics (Ai and Bhandari, 2021).
In the remainder of the paper, section 2 presents the model and section 3 establishes identification of the parameters. Section 4 discusses the empirical implementation, section 5 presents the results, and section 6 discusses welfare. Section 7 concludes. The supplemental appendix provides additional derivations and results.
2 Lifecycle model with tail income risk
The framework through which we study the extent of partial insurance to income shocks is a consumption lifecycle model with idiosyncratic income risk.
Household $i$ chooses lifetime consumption $\{C_{it}\}_{t=0}^{T}$ and savings $\{A_{it+1}\}_{t=0}^{T}$ to maximize its expected discounted utility over its lifecycle. The problem is formulated as

### v5

@@ Abstract
We propose a simple framework to measure consumption insurance against income shocks that accounts for higher-order moments of the income distribution. We derive a nonlinear consumption function in which the extent of insurance varies with both the sign and magnitude of shocks. Using recent PSID data, we estimate an asymmetric pass-through of bad versus good permanent shocks – 17% of a $3\sigma$ negative shock transmits to consumption versus 9% of an equal-sized positive shock – with greater pass-through as shocks worsen. These consumption transmission rates further vary by age, wealth, and along the income distribution. Our results align with survey responses to hypothetical events and suggest that tail risk matters substantially for consumption.
@@ 1 Introduction
How does consumption respond to income shocks? This question is central to understanding the welfare effects of shifts in the income distribution, i.e., the link between income and consumption inequality, how families cope after adverse events, government insurance design, and the dynamics of business cycles. A consistent empirical finding is that consumption is partially insured against income shocks, so income fluctuations translate into consumption fluctuations less than one-to-one (e.g., Blundell et al., 2008; Heathcote et al., 2014). [^1] The study of consumption insurance dates back to tests of the permanent income (Deaton and Paxson, 1994) and complete markets hypotheses (Cochrane, 1991; Attanasio and Davis, 1996; Hayashi et al., 1996; Brav et al., 2002). Jappelli and Pistaferri (2010) and Crawley and Theloudis (2024) review this literature. With a few exceptions discussed subsequently, most measures of consumption insurance do not distinguish between good and bad shocks, or between small and large ones. However, recent evidence shows that income risk has a long and fat left tail; that is, there is an asymmetrically large probability that households face some really bad income shocks. This paper proposes a simple framework to measure the consumption response to income shocks, precisely when the distribution of said shocks may exhibit long and fat tails.
@@
If most income shocks away from the mean are large and bad shocks, their effect on consumption might, intuitively, differ from that of small/moderate shocks. In other words, left skewness and excess kurtosis, which are statements about the third and fourth moments of the income distribution respectively, may matter for the pass-through of shocks into consumption. This is precisely what we investigate in this paper. Specifically, we measure the degree of consumption insurance to income shocks of varying persistence, accounting not only for second- but also higher-order moments of their distribution (i.e., tail shocks).
We offer three main contributions:
@@
First, we re-estimate BPP using recent PSID data over 1999–2019, targeting the original second moments of income and consumption, and compare the results with those based on imputed consumption from the contemporaneous CEX. The PSID previously lacked detailed consumption data, so BPP relied on an imputation from the CEX. However, such data are now internally available in the PSID. We subsequently use this data to study tail income risk, so it is natural to start with a ‘replication’ of BPP using the updated PSID. This exercise is insightful because of the different time period and the comparison with the CEX.
@@
Second, we introduce into BPP the third and fourth moments of the joint distribution of income and consumption, in addition to the second moments above. This is our first attempt to go beyond the workhorse income-consumption covariance and measure partial insurance targeting additional information in the tails of the distribution.
@@
Third, we generalize the method of measuring consumption insurance to explicitly permit a role for tail income risk. BPP log-linearize the policy rule of a lifecycle consumption/savings problem to obtain a consumption function that depends linearly on income shocks. Our first two contributions above revolve firmly around this linear framework. However, linearity may be a poor approximation to behavior when income is subject to tail shocks. We thus derive a higher-order (quadratic) consumption function that allows for nonlinear transmission of shocks. We characterize and empirically measure the consumption pass-through of income shocks, which now depends explicitly on the shocks’ magnitude and sign. While a nonlinear consumption function has been estimated before – most notably in Arellano et al. (2017), henceforth ABB – our simple framework offers key advantages: it separates the nature of income risk from the nature of the consumption response as distinct drivers of consumption growth along the income, age, and wealth distributions; it is far less data demanding (second moments of consumption and up to fourth moments of income suffice, with no need to observe the full consumption-income joint distribution or wealth); and it links identification and estimation transparently, aiding interpretation. We elaborate on these points below.
@@
Our core empirical exercises use income and consumption panel data from the PSID from 1999 to 2019. We focus on a representative sample of stable married couples that is similar to the sample in BPP. We show that unexplained income growth in this sample features large negative skewness and excess kurtosis, qualitatively and numerically close to the skewness and kurtosis in the administrative data of Guvenen et al. (2021). For the re-estimation of BPP in the recent years, we also use a comparable sample from the CEX.
We obtain four main findings:
@@
First, permanent and transitory shocks to disposable household income are negatively skewed, with skewness coefficients at $-0.95$ and $-1.31$, respectively, and highly leptokurtic, with kurtosis coefficients at $44.39$ and $49.23$, respectively. Left skewness implies that a negative shock is typically more unsettling than a positive shock, because it is farther away from the zero mean. Excess kurtosis implies that more shocks are concentrated either in the middle of the distribution or far out in the tails, relative to the shoulders. Taken together, large left skewness and kurtosis suggest that for most households the realizations of income shocks tend to be either small (i.e., close to the zero mean) or far out in the left tail. However, these features vary by age and by the household’s position in the income distribution.
@@
Second, in estimating BPP using the updated PSID in recent years, we find full insurance to transitory shocks and a small pass-through of permanent shocks at 0.15 versus 0.64 in BPP. A 10% permanent income cut thus reduces consumption by only 1.5%, vis-à-vis 6.4% in BPP. In other words, we find much larger partial insurance to permanent shocks in recent years than originally found in 1980–1992.222Using recent PSID data but a different method, ABB and Arellano et al. (2023) find a similarly low pass-through of permanent shocks. We discuss this further when we present our empirical results. We attribute about a third of the gap from BPP to the consumption imputation from the CEX that BPP use. The imputation brings in large amounts of measurement error, inflating the covariance between income and consumption growth. Another third of the gap arises because the modern PSID provides data biennially, versus annually previously, which tends to mute the transmission of higher frequency shocks. The last third reflects the different calendar times.
@@
Third, introducing the third and fourth moments of income and consumption growth into BPP’s linear consumption setting does not change our estimates of partial insurance. This is counterintuitive in light of the literature that highlights, in various settings, the large welfare costs of tail risk (e.g., Barro, 2009; Busch and Ludwig, 2024). One would thus expect that targeting higher-order moments reduces the degree of partial insurance, which is not what we find. This result is suggestive of a potential misspecification in the linear consumption function, which motivates our generalization of the methodology through the introduction of a higher-order consumption specification subsequently.
@@
Fourth, in estimating our quadratic consumption function, we find an asymmetry in the pass-through of large negative shocks versus positive or smaller ones. In this environment, the pass-through depends on the magnitude and sign of shocks; and targeting higher-order income moments enables the identification of the transmission parameters. We draw three main conclusions: 1.) Bad permanent shocks are more unsettling than equal-sized good ones because they have larger transmission rates into consumption – a three standard deviations negative shock (a 50% permanent income cut) has a pass-through of $0.17$ (17% of the shock transmits into consumption), while a positive shock of similar magnitude has a pass-through of $0.09$. 2.) The pass-through of bad shocks increases with the severity of the shock. Bad permanent shocks are thus double hurtful: not only are they more extreme than good ones (they lie farther out in the left tail), but they also transmit more strongly as their magnitude worsens. 3.) Transitory shocks are fully insured in the full sample, at least from a statistical point of view, although the point estimates of their transmission parameters suggest that large transitory windfalls shift consumption considerably, consistent with quasi-experimental evidence of large responses to transitory gains (e.g., Fagereng et al., 2021). Overall, the results are consistent with an environment in which households save a good income shock but dissave in the presence of a bad one. There is clearly a limit to how much one can dissave or borrow, so a large negative shock ends up affecting consumption by relatively more.
@@
The asymmetric transmission of bad versus good permanent shocks varies with age and the household’s position in the income distribution. While older households insure average permanent shocks better than younger ones, they respond more strongly to larger negative shocks. Moreover, bad permanent shocks in the lower tail of income risk cause more harm to the rich (households in the top of the income distribution) than to the poor (households in the bottom); the converse is true for good shocks.
How a shock from a given position in the distribution affects consumption depends on both the sign/size of the shock and the extent of its transmission. A key advantage of our moments-based approach is that it separates income risk and consumption transmission, allowing us to analyze each independently. Among the rich, the asymmetry is driven by an increasing severity of downside risk along the income distribution, rather than an increasing consumption transmission. Among the old, however, the asymmetry is because both the rate of transmission and the severity of downwards income risk increase with age.
The asymmetric transmission of bad versus good permanent shocks also varies with wealth and education. Transmission of a given average or tail shock decreases as we move from the least to the most wealthy/educated, so wealthier and highly educated households are better insured to all types of permanent shocks. However, the nature of income risk differs vastly across groups: the wealthy/educated face more severe downwards income risk, ultimately leading to larger consumption declines when bad shocks hit.
With left skewness and excess kurtosis, large negative shocks are more likely and more severe than positive ones, so households would rather remove such tail risk from their portfolio. A back-of-the-envelop calculation suggests that households facing large downwards risk would give up 4%-8% of lifetime consumption when income risk transmits linearly and rises to 6%-12%, almost up to 1/8 of lifetime consumption, in the presence of nonlinear transmission. There are several estimates of the welfare costs of income risk in the literature. Two settings in which the baseline environment concerns idiosyncratic tail income risk of similar nature and severity to ours are Busch and Ludwig (2024) and Guvenen et al. (2024). Agents in the former paper sacrifice between 0.5% and 14% of lifetime consumption to live in a world with Gaussian risk (a similar counterfactual to ours), while households in the latter sacrifice almost 22% of consumption to live in a world without risk (a more distant counterfactual from ours). [^3] These papers calculate welfare costs using full structural models in which the consumption responses depend on endogenously accumulated wealth. By contrast, we rely on approximate consumption rules, which take wealth as given. We show in appendix D that this is unlikely to be quantitatively important., [^4] Welfare cost estimates of aggregate risk include Barro (2009), who find that agents can give up about 20% of GDP each year to eliminate rare disasters, and Georgarakos et al. (2025), who find that households can sacrifice 5%-6% of lifetime consumption to eliminate business cycle fluctuations.
@@
There is a large literature in macro and labor economics that studies the transmission of income shocks into consumption, i.e., the link between income and consumption inequality. There are two distinct strands in this literature: a quasi-reduced form branch that measures the degree of partial insurance in the data without fully specifying the underlying insurance mechanisms, and a structural branch that fully models specific insurance channels. The present paper falls in the first branch.
@@
The seminal paper in the first branch, namely BPP, sparked a large volume of research. [^5] Hall and Mishkin (1982), Deaton and Paxson (1994) and Blundell and Preston (1998) characterize the empirical joint distribution of consumption and income, so they are early precursors in this literature. Blundell et al. (2016) measure partial insurance against wage risk, Theloudis (2021) allows for preference heterogeneity in partial insurance, and Commault (2022) allows consumption to respond to past transitory shocks and finds a strong response similar to what experimental studies find. Hryshko and Manovskii (2022) identify sets of households in the PSID with different degrees of insurance based on certain features of their incomes, and Crawley and Kuchler (2023) address neglected time aggregation in the PSID data. These papers restrict income shocks to transmit linearly into consumption. ABB and Arellano et al. (2023) generalize the income and consumption processes by letting income feature nonlinear persistence, which varies with the level of income and its shock, and nonlinear transmission that is more flexible than ours. Using recent PSID data and quantile methods, they measure the marginal propensity to consume from income along the income distribution. To quantify the pass-through of income shocks, they simulate the model and find asymmetries in the transmission of bad and good shocks, or between rich and poor households, similar to ours.
Our third contribution shares ABB’s goal of measuring nonlinear transmission, but differs in three key dimensions. First, ABB compare consumption growth across quantiles of income and income shocks, without distinguishing between the nature of income risk and that of the consumption response. For example, the rich appear to update consumption more strongly than the poor after a lower-tail shock, but this can reflect more severe downside risk or/and a stronger transmission independent of the shock. Our moments-based approach separates these channels, allowing us to analyze each independently. In this example, we find richer households respond less to a given shock (more insurance), but face worse downside risk, leading to larger overall consumption changes. Second, the quantile methods in ABB require observing the full consumption-income joint distribution, and wealth, while our approach requires only second moments of consumption and up to fourth moments of income. This makes our framework less data demanding, appealing both on its own and because, unlike income, we cannot benchmark higher-order consumption moments against population data. The price of our econometrically more accessible approach is that, unlike ABB, we maintain a linear income process (but subject to tail shocks). Yet, we find that nonlinear persistence is not crucial for the main consumption results, i.e., the asymmetry across different shocks and income levels. Third, our consumption insurance parameters are directly estimable in the data without simulations, and tightly linked to their identification. They are comparable to the existing literature on partial insurance and ‘portable’ as they describe transparently how shocks of different signs/magnitudes shift consumption.
@@
The asymmetric response to bad and good shocks is recently found in surveys that elicit responses to hypothetical events. Bunn et al. (2018) find that the marginal propensity to consume from bad shocks is larger than from good shocks; Christelis et al. (2019) and Fuster et al. (2020) find that responses to losses are larger than to gains, and the response increases with the severity of the loss. These papers consider observed hypothetical transitory shocks, while we document similar patterns in response to unobserved permanent shocks. [^7] Given our finding of heterogeneous responses to shocks, our paper is also broadly related to the literature that studies MPC heterogeneity (e.g., Lewis et al., 2022; Crawley and Kuchler, 2023).
The paper is finally related to the newer income dynamics literature that focuses on non-Gaussian aspects. Geweke and Keane (2000) and Bonhomme and Robin (2010) use the PSID to document, respectively, left skewness and excess kurtosis in earnings growth, while Guvenen et al. (2021) provide even richer evidence from administrative data. Guvenen et al. (2014) and Busch et al. (2018) study the cyclical nature of skewness in several countries. While these papers are mostly descriptive of the higher-order features of income, we further document the implications of these features for consumption. We thus contribute to a growing literature on the implications of tail risk, e.g., for growth (Barro, 2009), asset prices (Backus et al., 2011; Ghosh and Julliard, 2012; Constantinides and Ghosh, 2017), business cycles (McKay, 2017), and labor market dynamics (Ai and Bhandari, 2021).
In the rest of the paper, section 2 presents the income and consumption processes and section 3 discusses identification. Section 4 discusses the implementation, section 5 presents the results, and section 6 concludes. The appendix has additional derivations and results.
2 An organizing framework: a lifecycle model with tail income risk
The framework through which we organize the discussion of partial insurance to income shocks is a consumption lifecycle model with idiosyncratic income risk.
Household $i$ chooses a sequence of consumption $\{C_{it}\}_{t=0}^{T}$ and savings $\{A_{it+1}\}_{t=0}^{T}$ to maximize its expected discounted utility over its lifecycle. The problem is formulated as